OpenStudy (anonymous):

pete is 14 and his grandfather is 54. how many years ago was his grandfather 6 times as old as pete? step by step plz

OpenStudy (anonymous):

6(14-x)=54-x 84-6x=54-x 30=5x 6=x 6 years ago